OpenStudy (jamaljunior):

solve for x 2/7x+12/x=−40/7−1/6x

OpenStudy (kenrick143):

Its a quadratic equation.

OpenStudy (jamaljunior):

whats that mean

OpenStudy (kenrick143):

Theres two values of x

OpenStudy (jamaljunior):

so how do i start

OpenStudy (kenrick143):

Simplify both sides.

satellite73 (satellite73):

it is pretty much impossible to guess what you have written above if you use the equation tool below it would be clearer or attach the question if you can

OpenStudy (jamaljunior):

\[27x+12x=−407−16x\]

OpenStudy (jamaljunior):

ill post a picture i suppose cause thats not right i dont think

satellite73 (satellite73):

at the moment it looks like \[\frac{2}{7x}+\frac{12}{x}=-\frac{40}{7}-\frac{1}{6x}\] but i doubt that is right

OpenStudy (jamaljunior):

https://gyazo.com/bf4f63492bb607f0cb943f2371ca7127

satellite73 (satellite73):

ooh

OpenStudy (kenrick143):

So, you gave a different equation. Then you may start by getting rid of the fractions.

satellite73 (satellite73):

don't you hate fractions? multiply both sides by 42, cancelling as you go, and the fractions will be gone

OpenStudy (jamaljunior):

why multiply by 42?

OpenStudy (kenrick143):

To get rid of 7 and 6

satellite73 (satellite73):

because your denominators are 7,2 , 6 and 42 is the least common multiple of those three numbers